40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 26 03:57:06 2009, rob, Update, Locking, bad 
    Reply  Tue Oct 27 18:19:14 2009, rob, Update, Locking, bad MCF_issue.pngno_MCF_issue.png
       Reply  Thu Oct 29 21:51:07 2009, rob, Update, Locking, bad itsback.png
Message ID: 2141     Entry time: Mon Oct 26 03:57:06 2009     Reply to this: 2152
Author: rob 
Type: Update 
Category: Locking 
Subject: bad 

Lock acquisition has gone bad tonight. 

The initial stage works fine, up through handing off control of CARM to MCL.  However, when increasing the AO path (analog gain), there are large DC shifts in the C1:IOO-MC_F signal.  Eventually this causes the pockels cell in the FSS loop to saturate, and lock is lost. 

ELOG V3.1.3-